show step by step:

subtract ( 6x2 + 8x - 6) from ( 7x2 + 8x - 4)


1
Expert's answer
2021-04-15T07:52:41-0400

7x2+8x-4-(6x2+8x-6)=7x2+8x-4-6x2-8x+6=x2+2